OpenStudy (briellshaw):

Convert the fraction to a decimal. 5/6 A.0.56 B.0.83 C.5.6 D.8.3

OpenStudy (briellshaw):

@EclipsedStar @RennesmeeCullen

OpenStudy (briellshaw):

@TwilightRain1

OpenStudy (anonymous):

this should help... https://www.ltcconline.net/greenl/courses/187/c/ConvertingFractionsToDecimals.htm

OpenStudy (theweeknd16):

i think its B

OpenStudy (briellshaw):

How do you figure

OpenStudy (theweeknd16):

divide the numerator with the denominator.

OpenStudy (solomonzelman):

Start from diving \(\large\color{black}{ \displaystyle~~~ {\Huge \bf _\text{____} } }\) \(\large\color{black}{ \displaystyle 6 {\Large \bf|} 5 }\)

OpenStudy (briellshaw):

Thank you!!! I present you with a Medal, Testimonial, and 1 brand new fan

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle~~~ ^{\LARGE 0}_{\Huge \bf \text{____} } }\) \(\large\color{black}{ \displaystyle 6 {\Large \bf|} 5 }\)

OpenStudy (briellshaw):

@TwilightRain1

OpenStudy (anonymous):

it is b

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle~~~ _{\Huge \bf \text{____} }^{\Large 0.8} }\) \(\large\color{black}{ \displaystyle 6 {\Large \bf|} 50 }\) \(\large\color{black}{ \displaystyle ^{\bf~-} 48 }\) \(\large\color{black}{ \displaystyle ~~~^\text{___} }\) \(\large\color{black}{ \displaystyle ~~~~~~2 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle~~~ _{\Huge \bf \text{____} }^{\Large 0.~8~3} }\) \(\large\color{black}{ \displaystyle 6 {\Large \bf|} {\tiny~}50 }\) \(\large\color{black}{ \displaystyle ^{\bf~-} 48 }\) \(\large\color{black}{ \displaystyle ~~~^\text{___} }\) \(\large\color{black}{ \displaystyle ~~~~~~20 }\) \(\large\color{black}{ \displaystyle ~~~^{\bf~-} 18 }\) \(\large\color{black}{ \displaystyle ~~~~~~~^\text{___} }\) \(\large\color{black}{ \displaystyle ~~~~~~~~~2 }\)

OpenStudy (solomonzelman):

there is no latex for what I did just now using latex, and I am lagging... I will do better next putting it up next time.